OpenStudy (anonymous):

x+2/6=3/x-1 what doo youu doo first ?

OpenStudy (anonymous):

could u cross multiply

OpenStudy (anonymous):

multiply both sides by 6 & by (x-1): (x+2)(x-1)=3*6 x^2 +x-2=18 x^2 +x-20=0 solve it for x :)

OpenStudy (anonymous):

(x-4)*(x+5)=0 x=4 & x=-5

OpenStudy (anonymous):

thanks :3)